OpenStudy (christos):

Suppose that A = 1 L 0 0 L^2 3 2 1 1 where L belongs to R Find all the values of L in order for the Nullity A>0 (strictly positive).

OpenStudy (phi):

I would use row reduction to change the matrix to \[\left[\begin{matrix}1 & L&0 \\ 0 & L^2 & 3 \\ 0 & 1-2L &1\end{matrix}\right]\]

OpenStudy (phi):

(I add -2 times the first row to the last row) if we scale the last row by 3 \[ \left[\begin{matrix}1 & L&0 \\ 0 & L^2 & 3 \\ 0 & 3-6L &3\end{matrix}\right] \] we see we can can a repeated row (i.e. nullity >0 ) if \[ L^2 = 3-6L\] solve for L

OpenStudy (christos):

how is L ^2 equal to 3 - 6L @phi

OpenStudy (christos):

do you mean that I need to solve for L on that equation

OpenStudy (phi):

I am saying *if* L^2 equals 3-6L then the 2nd and 3rd rows will be duplicates

OpenStudy (christos):

*if* aha ok

OpenStudy (phi):

in this case, there are two (ugly) values for L that "work"